OpenStudy (anonymous):

Please Help!! Show work Rewrite y = x2 + 14x + 29 in general form. Rewrite y = 3x2 - 24x + 10 in general form. Solve for x: (x - 9)2 = 1 Solve for x: x2 + 24x + 90 = 0 Solve for x: 2x2 - 4x - 14 = 0 Create your own quadratic equation and demonstrate how it would be solved by graphing, factoring, the quadratic formula, and by completing the square.

OpenStudy (anonymous):

starting at solve 1. (x-9)2=1 2x-18=1 2x=19 x=9.5 2. x2+24x+90=0 26x+90=0 26x=-90 x=3\[\frac{ 6 }{ 13 }\]

OpenStudy (anonymous):

the fraction goes next to the three

OpenStudy (anonymous):

Wow Thank you!!!

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Those are the first 2?

OpenStudy (anonymous):

no the second two the first ones that you have to solve for x

OpenStudy (anonymous):

So how did you start at (x-9) =2?

OpenStudy (anonymous):

And so the final answer for number one is 3 6/13?

OpenStudy (anonymous):

@mathman806

OpenStudy (anonymous):

I'll have my sis help @sodukogirl13